OpenStudy (anonymous):

What are the solutions for x in the proportion x plus 6 all over 2 times x equals x over 2?

OpenStudy (swissgirl):

is it? \(\huge {x+6 \over 2x } = { x \over 2} \)

OpenStudy (anonymous):

yes :))

OpenStudy (swissgirl):

ok so we cross multiply meaninggg: \( (x+6)(2)=(2x)(x) Now we would have to distribute

OpenStudy (anonymous):

um so it would be 2x+2=2x^2?

OpenStudy (swissgirl):

yes it would be 2x+12=2x^2

OpenStudy (swissgirl):

Yes but then we need to bring over everything to one side \( 2x^2-2x+12 =0 \)

OpenStudy (anonymous):

okkkk

OpenStudy (swissgirl):

Its a parabola. Do you know what to do from here?

OpenStudy (anonymous):

um no how do you get two solutions for x if they are on the same side?

OpenStudy (swissgirl):

ok ill just give u the steps First you factor all that you can. In this particular equation we can factor out a 2 \( 2(x^2-x+6) \)

OpenStudy (swissgirl):

next we use that quadratic factoring We find two numbers that when you add them they equal -1 and when you multiply then it equals 6. So what do u think the 2 numbers are?

OpenStudy (anonymous):

-3 and 2?

OpenStudy (swissgirl):

yaaaaaaaaaaaaaaaaa

OpenStudy (anonymous):

thankkkk youu soo much!! can you help me with 1 more pleasee??? :)

OpenStudy (swissgirl):

so now we replace the -x with 2x-3x which is in essence -x \( 2( x^2+2x-3x-6)=0 \) \( 2((x^2+2x)-(3x+6))=0 \) \( 2(x(x+2)-3(x+2)) =0\) \(2((x-3)(x+2)) =0 \) \( (x-3)=0 ,and, (x+2)=0 \) x=3 and x=-2

OpenStudy (swissgirl):

So your x equals 3 and -2 NOT -3 and 2

OpenStudy (swissgirl):

I hope u come back on and see this, I was just busy with smth else and didnt have time to respond